OpenStudy (anonymous):

what's the general term of S=4,12,36,...,708588

OpenStudy (anonymous):

tn=ar^(n-1) n>=1

OpenStudy (anonymous):

tn=ar^(n-1) ; n>=1

OpenStudy (anonymous):

but how do I know the values?

OpenStudy (nikvist):

a=4,r=3